# A Secure Quantum Key Distribution Protocol Using Two-Particle Transmission

### v2

@@ 1 Introduction
Contemporary QKD primarily relies on non-orthogonal states for security. However, the adoption of orthogonal states in cryptographic protocols emerged later, with the inception of the pioneering protocol [25]. This groundbreaking approach introduced the concept of sending states with controlled time delays, making it nearly impossible for eavesdroppers to intercept an entire state without detection. Another method, involving order rearrangement, facilitates the implementation of QKD protocols based on orthogonal states, as demonstrated by various documented protocols [26, 27, 28, 29]. Experimental validation has recently been conducted in the case of quantum cryptography based on orthogonal states [30, 31]. While non-orthogonal state encoding is prevalent, orthogonal state encoding offers potential benefits, such as reduced quantum operation requirements. Understanding the theoretical application of orthogonal states for coding is invaluable due to their innate ability to be distinguished without errors.
@@ 2 Protocol
where $A$ and $B$ represent the states of particles $A$ and $B$, respectively. The bases given by Eq.2 are not complete, indicating that $\sum_{i=1}^{5}|\psi_{i}\rangle\langle\psi_{i}|\neq\mathbb{I}$, where $\mathbb{I}$ represents the identity matrix. However, it’s important to note that these bases are orthogonal to each other, meaning that $\langle\psi_{i}|\psi_{j}\rangle=0$ for all $i\neq j$ within the set $\{1,2,...,5\}$. Additionally, they exhibit a degree of nonlocality even in the absence of entanglement. Notably, these bases are indistinguishable under the framework of LOCC even in scenarios with infinite resources and infinite number of iterations [37, 38, 46, 47, 48, 49].
@@ 4 Conclusion:
Our investigation into utilizing $3\times 3$ tile UPBs for establishing secure quantum keys highlights their indistinguishability under the LOCC paradigm. While our protocol restricts eavesdroppers’ access to two-particle states, a time gap between transmissions limits adversaries to LOCC tools for eavesdropping. We acknowledge the potential existence of better protocols for utilizing UPBs in QKD and generalizing them for $d\times d$ dimensions. To extend our protocol, one may follow the same steps and formulate a valid quantum measurement. Let $|\psi_{1}\rangle,\ldots,|\psi_{l}\rangle$ form the UPB in $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ [57, 58], where $\sum_{i=1}^{l}|\psi_{1}\rangle\langle\psi_{i}|\neq\mathbb{I}$ and $|\psi_{l}\rangle$ is the stopper basis. Construct complete measurement bases by first forming orthogonal bases $|h_{1}\rangle,\ldots,|h_{l-1}\rangle$ orthogonal to each state of $|\psi_{1}\rangle,\ldots,|\psi_{l-1}\rangle$, with $\langle h_{k}|\psi_{i}\rangle=0$ for $i=1,\dots,l-1$ and $k=1,\ldots,l-1$. Then, formulate:
$|\psi_{l+1}\rangle$ $=$ $\alpha_{l+1}\Big{(}|h_{1}\rangle-\sum_{i=1}^{l}\langle\psi_{i}|h_{1}\rangle|\psi_{i}\rangle\Big{)}$
$|\psi_{k}\rangle$ $=$ $\alpha_{k}\Big{(}|h_{k-l}\rangle-\sum_{i=1}^{l}\langle\psi_{i}|h_{k-l}\rangle|\psi_{i}\rangle-\sum_{j=l+1}^{d^{2}-1}\langle\psi_{j}|h_{k-l}\rangle|\psi_{j}\rangle\Big{)}\;\;\;\;\text{for}\;\;k=l+2,\ldots,d^{2}$ (6)
such that $\sum_{i=1}^{d^{2}}|\psi_{i}\rangle\langle\psi_{i}|=\mathbb{I}$. Here, $\alpha_{k}$ is the normalization constant of the state $|\psi_{k}\rangle$. Using our protocol, one can distribute the quantum key between distant parties using the $d\times d$ UPBs. Exploring the optimal intercept resend attack probability would be a compelling avenue for future research.

### v3

@@ 1 Introduction
Contemporary QKD primarily relies on non-orthogonal states for security. However, the adoption of orthogonal states in cryptographic protocols emerged later, with the inception of the pioneering protocol [25]. This groundbreaking approach introduced the concept of sending states with controlled time delays, making it nearly impossible for eavesdroppers to intercept an entire state without detection. Moreover, several other studies facilitate the implementation of QKD protocols using orthogonal states, as evidenced by various documented protocols [26, 27, 28, 29]. Experimental validation has recently been conducted in the case of quantum cryptography based on orthogonal states [30, 31]. While non-orthogonal state encoding is prevalent, orthogonal state encoding offers potential benefits, such as reduced quantum operation requirements. Understanding the theoretical application of orthogonal states for coding is invaluable due to their innate ability to be distinguished without errors.
@@ 2 Protocol
where $A$ and $B$ represent the states of particles $A$ and $B$, respectively. The bases given by Eq.2 are not complete, indicating that $\sum_{i=1}^{5}|\psi_{i}\rangle\langle\psi_{i}|\neq\mathbb{I}$, where $\mathbb{I}$ represents the identity matrix. However, it’s important to note that these bases are orthogonal to each other, meaning that $\langle\psi_{i}|\psi_{j}\rangle=0$ for all $i\neq j$ within the set $\{1,2,...,5\}$. Additionally, they exhibit a degree of nonlocality even in the absence of entanglement. Notably, these bases are indistinguishable under the framework of LOCC [37, 38, 46, 47, 48, 49].
@@ 4 Conclusion:
Our investigation into utilizing $3\times 3$ tile UPBs for establishing secure quantum keys highlights their indistinguishability under the LOCC paradigm. While our protocol restricts eavesdroppers’ access to two-particle states, a time gap between transmissions limits adversaries to LOCC tools for eavesdropping. We acknowledge the potential existence of better protocols for utilizing UPBs in QKD and generalizing them for $d\times d$ dimensions. To extend our protocol, one may follow the same steps and formulate a quantum measurement. Let $|\psi_{1}\rangle,\ldots,|\psi_{l}\rangle$ form the UPB in $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ [57, 58], where $\sum_{i=1}^{l}|\psi_{i}\rangle\langle\psi_{i}|\neq\mathbb{I}$ and $|\psi_{l}\rangle$ is the stopper basis. Construct complete measurement bases by first forming orthogonal bases $|h_{l+1}\rangle,\ldots,|h_{d^{2}}\rangle$ orthogonal to each basis of $|\psi_{1}\rangle,\ldots,|\psi_{l-1}\rangle$, with $\langle h_{k}|\psi_{i}\rangle=0$ for $i=1,\dots,l-1$ and $k=l+1,\ldots,d^{2}$ and then, formulate:
$|\psi_{l+1}\rangle$ $=$ $\alpha_{l+1}\Big{(}|h_{l+1}\rangle-\sum_{i=1}^{l}\langle\psi_{i}|h_{l+1}\rangle|\psi_{i}\rangle\Big{)}$
$|\psi_{k}\rangle$ $=$ $\alpha_{k}\Big{(}|h_{k}\rangle-\sum_{i=1}^{l}\langle\psi_{i}|h_{k}\rangle|\psi_{i}\rangle-\sum_{j=l+1}^{k-1}\langle\psi_{j}|h_{k}\rangle|\psi_{j}\rangle\Big{)}\;\;\;\;\text{for}\;\;k=l+2,\ldots,d^{2}$ (6)
such that $\sum_{i=1}^{d^{2}}|\psi_{i}\rangle\langle\psi_{i}|=\mathbb{I}$. Here, $\alpha_{k}$ is the normalization constant of the state $|\psi_{k}\rangle$. Let’s say, if the number of UPB in a $d\times d$ system is $d^{2}-2d+1$, then one has to form $12d-$ 1 number of entangled bases to construct a complete set of measurement bases [57]. Using our protocol, one can distribute the quantum key between distant parties using the $d\times d$ UPBs. Exploring the optimal intercept resend attack probability would be a compelling avenue for future research.